On January 1, 2019, Joan Campbell borrows $20,000 from Susan Rone and agrees to repay this amount in payments of $4,000 a year until the debt is paid in full. Payments are to be of an equal amount and are to include interest at 12% on the unpaid balance of principal at the beginning of each period. Assuming that the first payment is to be made on January 1, 2020, determine the number of payments of $4,000 each to be made and the amount of the final payment.


What is the future value on December 31, 2026, of 7 annual cash flows of $10,000 with the first cash payment made on December 31, 2019, and interest at 12% being compounded annually?


1
Expert's answer
2022-03-07T10:44:13-0500

1.) Joan Campbell repayment schedule


"Year 1=4000+(0.12\times20000)=6,400"


"Year 2=4000+(0.12\times16000)=5920"


"Year 3=4000+(0.12\times12000)=5440"


"Year 4=4000+(0.12\times8000)=4960"


"Year 5=4000+(0.12\times4000)=4480"


Number of payments is 5 and final payment is $4,480


2.) Future Value.

"FV=A\times\frac{(1+r)^{7}-1}{r}"


"FV=10,000\times\frac{(1+0.12)^{7}-1}{0.12}=100,890.12"


Future Value= $100,890.12